OpenStudy (anonymous):

For the equation given below, evaluate y' at the point (-1,1). (2 x - y)^4 + 2 y^3 = 83

OpenStudy (anonymous):

start with \[4(2x-y)^3(2+y')+6y^2y'=0\] replace \(x\) by \(-1\) , \(y\) by \(1\) and solve for \(y'\)

OpenStudy (anonymous):

oops typo there \[4(2x-y)^3(2-y')+6y^2y'=0\]

OpenStudy (anonymous):

-108(2-y')+6y'=0?